OpenStudy (anonymous):

Factor 4x^2 - 20x +25

OpenStudy (anonymous):

You could use the quadratic formula to help you factor it

OpenStudy (hba):

@ChmE Are you sure ?

OpenStudy (anonymous):

it will get you the 2 solutions then you could just setup the ( )( ) form from there.

OpenStudy (hba):

@ChmE Sorry to say you are wrong.

OpenStudy (anonymous):

a=4 b=-20 c=25

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

when i worked it out it is (4x-5)(x-5) but checking it, it doesnt work

OpenStudy (hba):

@ChmE Do you know what the standard form should be ? It should be \[ax^2+bx+x=0\]

OpenStudy (hba):

Not \[ax^2+bx+x\]

OpenStudy (anonymous):

\[20 \pm \sqrt{20^2-4(4)(25)} \over 2(4)\]\[20 \pm \sqrt{400-400} \over 8\]\[{20 \over 8}={5 \over 2}\]This will be the only solution, but will happen twice (x-(5/2))(x-(5/2)) times 2 to get rid of the fraction (2x-5)(2x-5) This could have been done by factoring but you could use the quadratic formula to help you if you couldn't figure it out

OpenStudy (hba):

Doing it the wrong way does not help when you have to show the solution to the teacher.

OpenStudy (anonymous):

Factoring: 4x^2 - 20x + 25 Split the middle term 4x^2 - 10x - 10x + 25 Group the left 2 terms and the right two terms, and factor them 2x(2x - 5) + -5(2x-5) factor out the (2x-5) (2x-5)(2x-5)

OpenStudy (anonymous):

I don't know what your problem is @hba , but either get a purple border around your picture or shut up

OpenStudy (hba):

I am just trying to show him the right way,stop being rude.

OpenStudy (anonymous):

There is no "right", there are two possible ways, whatever is easier for that person is the right way for them.

OpenStudy (hba):

@ChmE But that is not the correct way, you are not understanding that.